OpenStudy (anonymous):

Absolute Value Equations and Inequalities? Please help! 1. Which conjunction or disjunction is equivalent to the open sentence |a – 8| ≤ 3. A. a – 8 ≤ 3 and a – 8 ≥ –3 B. a – 8 ≥ 3 and a – 8 ≤ –3 C. a – 8 ≤ 3 or a – 8 ≥ –3 D. a – 8 ≥ 3 or a – 8 ≤ –3

OpenStudy (amoodarya):

you know |8-a|=|a-8| that is better now you may know |x|<k --> -k<x<k so |a-8|<=3 -3<=a-8<=3 8-3<=a<=3+8 can you go on?

OpenStudy (anonymous):

\[\left| x-a \right|\le b,-b \le x-a \le b\] \[\left| x-a \right|\ge b,x-a \le -b and x-a \ge b\]